• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Meziani

    Log entry time 00:03:31 on July 8,2007

    Entry number 207974

    keyword=Shift Summary

    16:00 Shift start
    ps1 = set to 30
    Current at 30 uA
    16:05-16:37 run # 4403 300K good
    Current raised to 60 uA
    16:37-16:50 run # 4404 300K
    16:51-17:09 run # 4405 300K
    17:10-17:24 run # 4406 300K
    17:24-17:41 run # 4407 300K
    17:41-17:56 run # 4408

    Start a density study vs current

    17:56-18:01 run 4409 at 65 uA

    18:02-18:05 run 4410 at 55 uA but beam tripped
    18:07-18:12 run 4411 at 55 uA
    18:15-18:22 run 4412 at 45 uA
    18:24-18:32 run 4413 at 35 uA: Current mainly at 36 + a sudden rise to 38 for a 10sec.
    18:38-18:43 run 4414 at 25 uA
    18:45-18:48 run 4415 at 15 uA
    18:50-18:52 run 4416 at 5uA tripped
    18:52-18:55 run 4417 at 5 uA

    Density study finished, start production runs again
    19:00-19:45 run # 4418 1M events 4He production
    19:45-20:30 run# 4419 1M #
    20:30-21:10 run# 4410 1M #

    Next is a 20cm Dummy target run at 40 uA
    21:21-21:49 run# 4421 420 K, Dummy 20cm

    Back to production runs at 65 uA
    21:57-22:36 run # 4422 1M events 4He
    22:37-23:16 run # 4423 1M #
    23:16-23:57 run # 4424 1M #
    Asked for controlled access for David Meekins to go in.
    24:00 Shift ends